OpenStudy (anonymous):

find the rate of change of C with respect to P: C(p)=1100p/140-p

OpenStudy (da_scienceman):

Just compute the derivative!!

OpenStudy (anonymous):

C(p)=f(x)/g(x) C'(p)=[g(x)f'(x)-f(x)g'(x)] / g(x)² This is what you'll find if you carry out the operation based on the definition of a derivative... the quotient rule (or something like that).

OpenStudy (da_scienceman):

yes and f(x) = 100p, g(x) = 140-p, substitute like u said and simplify!